# Complementary Learning for Real-World Model Failure Detection

### v1

@@ IV-A Model Failure Classification
Figure 7: Discrepancy Detection. Given mostly normal data in the KITTI Odometry dataset, the two streams agree on 95 % of the data (green, blue). In light of anomalies, as provided by the CODA datasets, many more discrepancies (red, yellow) are detected.
Next, we show scenarios where the self-supervised stream showed model failures, detected by correct predictions of the supervised stream. Figure 8 shows representative scenes of this category. Scene 1 contains two distant pedestrians walking, wrongly classified as static by the self-supervised stream. In scene 2, a parked car is misclassified as dynamic by the self-supervised stream. Scenes 3, 4, and 5 show walking pedestrians or moving cars incorrectly classified as dynamic. These cases demonstrate that our approach enables the detection of challenging temporal scenarios. Often, models under test are fine-tuned variants of models trained in a self-supervised fashion. This way, additional training data can be collected for improved pre-training. Here, no labels are required. The self-supervised stream classifies an above-average number of objects as dynamic when the ego-vehicle turns or goes over a speed bump. An example is shown in Figure 8, where in scene 6, the vehicle turns, and in scene 8, it drives over a speed bump. Another weak point is fast oncoming vehicles on highways, which are often classified as static, as can be seen in scene 7. Finally, a common weakness is small clusters on the right or left edge that are incorrectly classified as dynamic, as in scene 9, where points of a window are classified as dynamic.
Finally, also cases occur where the models are incorrectly consistent. For this category, the two streams agree, but the label is incorrect in both cases. We show examples in Figure 10. Here, the left scene shows two walking pedestrians that are incorrectly classified as static and the right scene shows a parked car classified as dynamic.
@@
Figure 8: Self-Supervised Model Failures. Based on our visual inspection tool, a human oracle can analyze contradicting model outputs and classify failure cases. In these exemplary images, we show model failures of the self-supervised model, which can be detected due to contradicting outputs of the supervised stream. Adapted from [48].
IV-B Sensitivity towards Anomalies
Generally, we are interested in scenarios where models fail. It is well known that perception models often struggle most with anomalies in the environment. Thus, we also examine the sensitivity of our approach towards object-level anomalies in the environment around the ego vehicle, as these are most often examined [65, 66].
Evaluation Data. While our goal in the first part of the evaluation was to minimize the domain gap between training and evaluation data, there are no labeled anomalies in the original KITTI dataset [64]. Thus, we utilized data from CODA [67], the only real-world dataset that provides lidar data and includes anomalies [68]. The CODA dataset provides anomaly labels for objects based on three existing data sets: KITTI [64], ONCE [69], and nuScenes [70].
@@ IV-B Sensitivity towards Anomalies
For the CODA-KITTI split, the authors manually reviewed all misc labels available in the ground truth and relabeled some as anomalies according to a labeling policy. Since we trained our models on KITTI-360, this enables us to quantitatively examine our approach with only a small domain gap. For CODA-nuScenes, the authors similarly adopted available annotations in a manual process. Finally, for CODA-ONCE, they deployed an automated anomaly detection approach, making this subset the most relevant. CODA includes 1,500 scenes with a total of 5,937 anomaly instances. Among those, 4,746 belong to the superclass traffic_facility, followed by 929 vehicle and 197 obstruction instances. With 396, most vehicle instances can be found in CODA-KITTI. In Figure 7, we show the outputs of our approach on the CODA subsets, also in comparison to the outputs of the KITTI dataset. We can observe many more detected discrepancies, which is in line with the much higher number of anomalies, but the subsets reveal strongly varying behavior patterns.
For further qualitative analysis, however, labeled anomalies in 3D are required. The original CODA dataset provides anomaly labels only in the form of 2D bounding boxes in images. Therefore, we present LidarCODA, a dataset based on the CODA dataset [67] for evaluation. Based on a frustum-based filter, subsequent clustering, and manual inspection, we transferred the original 2D labels from image space into refined, point-wise 3D labels that go beyond the coarse characteristic of the provided bounding boxes. More details can be found in [71]. LidarCODA is the first real-world anomaly dataset with annotated lidar data, as shown in Figure 9. Here, also the different lidar systems utilized become clearly visible. Due to the sparse point cloud of nuScenes, many small or distant labeled anomalies in the image space are only covered by a few or no lidar points.
Dataset #Frames mIoU $\uparrow$ mP $\uparrow$ mR $\uparrow$ mF1 $\uparrow$
@@
TABLE I: LidarCODA Subsets. Evaluation of our approach on LidarCODA and its three subsets. Adapted from [71].
Anomalies. With LidarCODA, we provide ground truth for object-level anomalies [1] in lidar data. Since our approach is not specific to object-level anomalies but shows its strength in underrepresented or generally challenging or atypical scenarios, which can also include scene- or scenario-level anomalies, we are primarily interested in the sensitivity of our approach towards these specific anomalies. This is different from works in anomaly detection, which aim to detect as many anomalies as possible. As our approach assigns distinct labels per point, we follow the evaluation protocols of semantic segmentation tasks rather than anomaly detection tasks, as here an anomaly score per point is required. First, we need to better understand the suitability of the subsets of LidarCODA due to introduced domain shifts, either due to new environments or due to new sensor setups. Table I shows the evaluation results on the individual subsets. Here, we evaluated all points of the lidar point cloud, even if our approach did not label individual points, e.g., because they were filtered during pre-processing. Such cases were counted as false negatives if an anomaly was missed. This way, we fairly incorporate the limits of our approach and do not create any requirements on the analyzed point clouds. The results clearly show that our approach struggles with the nuScenes subset, which is primarily due to the large domain shift w.r.t to the sensor setup. The subsets ONCE and KITTI, however, show more promising results, as here, our approach is capable of detecting anomalies. This allows for further investigation. As nuScenes is by far the smallest subset, its impact on the evaluation is rather limited.
Next, we investigate the sensitivity towards the superclasses provided by CODA. Here, we only evaluated points that have labels assigned by both our approach and the ground truth. As shown in Table II, our approach shows different levels of sensitivity given different types of anomalies. While animals were not detected at all, our approach is more sensitive to cyclists and objects of the misc class. The misc class consists of objects that are “unrecognizable or difficult to categorize” [67]. The results are difficult to interpret, however, as CODA defines an anomaly as an object that “blocks or is about to block a potential path of the selfdriving vehicle” [67] and/or “does not belong to any of the common classes of autonomous driving benchmarks” [67]. This risk-aware definition is not always in line with the methodology of our approach, where objects that block the path in front of the ego vehicle are not necessarily hard to segment or predict. Regardless, we can conclude that our methodology shows a heightened sensitivity towards cyclists who are sometimes difficult to recognize or predict and anomalies that are generally difficult to categorize.
@@ V Conclusion
Superclass #Instances mIoU $\uparrow$ mP $\uparrow$ mR $\uparrow$ mF1 $\uparrow$
@@
TABLE II: CODA Superclasses. Evaluation of our approach on seven superclasses. Adapted from [71].
In this work, we have presented complementary learning to detect real-world model failures. We leverage complementary training paradigms to detect contradicting outputs on the same task. This way, we can detect the failure modes of a model in the real-world, far beyond the limited scope of a test split. We demonstrate our approach with a motion segmentation of point clouds for autonomous driving, where we leverage a supervised stream for semantic motion labels and a self-supervised stream for predictive motion labels. By inspecting over 20,000 frames, we showed that our approach exposes model failures in repeating scenarios, which allows for the classification of model failures. In addition, we provide LidarCODA, the first anomaly detection benchmark with labeled point clouds to perform a quantitative analysis. We could show that our approach shows an increased sensitivity of often hard-to-detect and hard-to-predict bicycles as well as hard-to-categorize objects. In the future, we will be interested in the effects of including challenging data points into re-training our models, and potentially fine-tuning our self-supervised stream with the supervised one.
Limitations. When both streams are wrong, model failures go undetected, as shown in Figure 10. This behavior is known and unavoidable [21, 33]. Thus, our approach should not be understood as a stand-alone solution but can be seen as one of many triggers to detect challenging scenarios. In addition, our self-supervised stream depends on a clustering strategy, which can lead to faulty clusters and misclassifications.
@@ Acknowledgment
This work results partly from the jbDATA project supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), grant number 19A23003H.

### v2

@@ IV-A Regular Scenarios
Figure 6: Discrepancy Detection. The charts show the distribution of the four possible outcomes of the discrepancy detection module for four different datasets. Given mostly normal data in the KITTI Odometry dataset, the two streams agree on 95 % of the data (green, blue). For outlier scenarios, as provided by LidarCODA, many more discrepancies (red, yellow) are detected.
First, we discuss model failures of the supervised stream, which is the model under test in most cases. Here, the two streams contradict each other. We show representative examples in Fig. 7. Scene 1 shows a turning car and two moving bicyclists, where one bicyclist is wrongly labeled as static by the supervised stream. Scene 2 contains two walking pedestrians that are wrongly classified as static by the supervised stream. Scene 3 shows a parked car misclassified as dynamic by the supervised stream. Scenes 4 and 5 show a car moving slowly and a car moving backward, respectively. These cases demonstrate effectively that our approach enables the detection of regular but challenging scenarios that lead to model failures. Such model failures remain undetected in small evaluation datasets. We find various weak points in each stream, characterized by repeated occurrence. Specifically, the supervised model under test has weaknesses in distinguishing between dynamic and static objects in specific situations, e.g., at red lights or when a car is parked directly in front of the ego vehicle. Examples of such situations are given in scenes 6 and 7 of Fig. 7.
@@
Figure 7: Supervised Model Failures. These exemplary images show model failures of the supervised stream, which can be detected due to contradicting outputs of the self-supervised model. Adapted from [48].
Next, we analyze scenarios where self-supervised model failures occur, detected by correct predictions of the supervised stream. Fig. 8 shows representative scenes. Scene 1 contains two distant pedestrians walking, wrongly classified as static by the self-supervised stream. In scene 2, a parked car is misclassified as static by the self-supervised stream. Scenes 3, 4, and 5 show walking pedestrians or moving cars incorrectly classified as dynamic. These cases demonstrate that our approach enables the detection of challenging temporal scenarios. The self-supervised stream classifies an above-average number of objects as dynamic when the ego-vehicle turns or goes over speed bumps. An example is shown in Fig. 8, where in scene 6, the vehicle turns, and in scene 8, it drives over a speed bump. Another weak point is fast oncoming vehicles on highways, often classified as static, as seen in scene 7. Finally, a common weakness is small clusters on the side, which are incorrectly classified as dynamic, as in scene 9, where a window is classified as dynamic.
In rare cases, both models are incorrectly consistent, i.e., both streams agree, but the label is incorrect in both cases. We show examples in Fig. 9. Here, the left scene shows two walking pedestrians that are incorrectly classified as static, and the right scene shows a parked car that is classified as dynamic.
Figure 8: Self-Supervised Model Failures. These exemplary images show model failures of the self-supervised model, which can be detected due to contradicting outputs of the supervised stream. Adapted from [48].
IV-B Outlier Scenarios
Scenarios with anomalies or outliers are known to lead to model failures [70]. Based on state-of-the-art datasets, evaluating lidar-based anomaly detection models has been challenging. Evaluation datasets are either unavailable [71, 72] or utilize known classes but exclude them from training data [73]. Thus, we first introduce LidarCODA, a real-world anomaly dataset with labeled anomalies in lidar space. Subsequently, we quantitatively evaluate how sensitive our approach is to anomalies in the environment.
Evaluation Data. For the evaluation, we utilized data from the CODA dataset [69], the only publicly available real-world dataset that provides lidar data and includes anomalies [74]. The CODA dataset provides anomaly labels for objects based on three existing datasets: KITTI [75], ONCE [5], and nuScenes [1]. CODA defines an anomaly as an object that “blocks or is about to block a potential path of the self-driving vehicle” [69] and/or “does not belong to any of the common classes of autonomous driving benchmarks” [69]. While the first risk-aware definition is not always in line with the methodology of our approach, where objects that block the path in front of the ego vehicle are not necessarily hard to segment or predict, the second one is well-suited. Novel classes are often more challenging for supervised methods compared to self-supervised approaches.
For the CODA-KITTI split, the authors of CODA manually reviewed all misc labels available in the ground truth and relabeled some as anomalies according to a labeling policy. This enables us to examine our approach quantitatively with only a small domain gap. For CODA-nuScenes, the authors similarly adopted available annotations in a manual process. Finally, for CODA-ONCE, they deployed an automated anomaly detection approach, making this subset the most relevant. CODA includes 1,500 scenes with a total of 5,937 anomaly instances. Among those, 4,746 belong to the superclass traffic_facility, followed by 929 vehicle and 197 obstruction instances. With 396, most vehicle instances can be found in CODA-KITTI.
The CODA dataset provides anomaly labels only in the form of 2D bounding boxes in image space. However, point-wise labels in 3D lidar space are necessary to utilize CODA for our approach. Therefore, we present LidarCODA, a dataset based on the CODA dataset [69] for evaluation. Based on a frustum-based filter, subsequent clustering, and manual inspection, we transfer the original 2D labels from image space into refined, point-wise 3D labels that go beyond the coarse characteristic of the provided bounding boxes. More details can be found in [76]. LidarCODA is the first real-world anomaly dataset for object-level anomalies [63] with annotated lidar data, as shown in Fig. 10. Here, the different lidar systems utilized also become clearly visible. Due to the sparse point cloud of nuScenes, many small or distant labeled anomalies in the image space are only covered by a few or no lidar points.
@@ IV-B Outlier Scenarios
TABLE I: LidarCODA Subsets. Evaluation of our approach on LidarCODA and its three subsets. Best results bold and second-best underlined. Adapted from [76].
Dataset #Frames mIoU $\uparrow$ AP $\uparrow$ AR $\uparrow$ F1 $\uparrow$
@@
Evaluation. To quantitatively evaluate the correlation between model disagreements and anomalies present in the environment, we utilize the standard metrics mean Intersection over Union (mIoU), Average Precision (AP), Average Recall (AR), and F1 score, as shown in Tables I and II. For a fair evaluation, we consider all points of the lidar point cloud, even if our approach did not label individual points, e.g., because they were filtered out during pre-processing. Such cases were counted as false negatives if an anomaly was missed.
To better understand the suitability of LidarCODA under introduced domain shifts, either due to new environments or due to new sensor setups, we perform experiments on the individual subsets, as shown in Table I. The results clearly show that our approach struggles with the nuScenes subset, which is primarily due to the large domain shift with respect to the sensor setup. Our approach is more sensitive towards anomalies for the subsets ONCE and KITTI. This is reflected in Fig. 6, where the CODA subsets also show much higher detection rates of model failures compared to the analysis with regular scenarios. This aligns with the much higher number of anomalies, even though the subsets reveal strongly varying behavior patterns.
Next, we investigate the sensitivity of our approach towards the superclasses provided in CODA by evaluating points that have labels assigned by both our approach and the ground truth. As shown in Table II, our approach shows different levels of sensitivity given different types of anomalies, being most sensitive to cyclists and objects of the misc class. The model performs worst on the class animal, which is difficult to interpret given the small number of only five instances. The misc class contains objects that are “unrecognizable or difficult to categorize” [69]. These results align well with our approach, where cyclist instances, which are hard to predict by the self-supervised stream, and misc instances, which are rare and thus hard to classify by the supervised stream, lead to model disagreements, as the complementary model does not display similar weaknesses in both cases.
@@ V Conclusion
TABLE II: CODA Superclasses. Evaluation of our approach on seven superclasses. Best results bold and second-best underlined. Adapted from [76].
Superclass #Instances mIoU $\uparrow$ AP $\uparrow$ AR $\uparrow$ F1 $\uparrow$
@@
In this work, we have presented a label-free approach to detect model failures for the segmentation of point clouds, focusing on the classes static and dynamic. We leverage complementary learning paradigms to detect contradicting outputs on the same task, consisting of a supervised stream for semantic motion labels and a self-supervised stream for predictive motion labels. This way, we detect model failures far beyond the limited scope of a small evaluation dataset. We first inspect model failures in regular scenarios. By manually analyzing over 20,000 frames qualitatively, we detect model failures in seemingly normal scenarios and are able to categorize frequently occurring cases. In regular scenarios, our method categorizes 95 % of the data as typical, which makes human analysis of the remaining 5 % feasible even for larger datasets.
Second, we analyzed the sensitivity of our approach towards scenarios with anomalies, as defined by the CODA dataset, and showed a moderate sensitivity. In order to quantitatively examine our approach, we introduced LidarCODA, the first real-world dataset with labeled anomalies in lidar space. We demonstrated that our approach shows an increased sensitivity to often hard-to-detect and hard-to-predict bicycles as well as hard-to-categorize objects.
Our approach effectively unveils model failures far beyond those that can be detected with small evaluation datasets. This leads to an increased understanding of the model performance in large-scale deployments, leveraging abundantly available unlabeled data. Model failures detected by our approach can be utilized to collect additional training data representing both static and temporally challenging scenarios.
Limitations. When both streams are wrong, model failures go undetected. This behavior is known and unavoidable [18, 31] and can be mitigated by deploying multiple approaches or triggers to detect challenging scenarios [9]. In addition, our self-supervised stream depends on a clustering strategy, which can lead to faulty clusters and misclassification.
Future Work. The proposed complementary learning approach can be applied to the detection of model failures in different settings that can be addressed with different training paradigms, such as full panoptic segmentation beyond the utilized static and dynamic classes [77], lane detection [78], or drivable-area segmentation [79].
@@ Acknowledgment
This work results partly from the just better DATA project supported by the German Federal Ministry for Economic Affairs and Climate Action (BMWK), grant number 19A23003H.